OpenStudy (anonymous):

if the radius of a circle is increased by 10% then the area is increased by?

OpenStudy (anonymous):

pi*r^2

OpenStudy (anonymous):

will the increase in area be 0.03% @.Sam. ?

OpenStudy (amriju):

take the initial radius to be r...then enlarged radius is r+r/10=11r/10...now area=\[\Pi*radius ^{2}\] .....so u have initial area as 3.14r^2...and final as 3.14*121r^2/100....so increase is 21/100*pi*r^2...thats 21 percent increase

OpenStudy (anonymous):

how it's 11r? @amriju i didnt get u :(

OpenStudy (amriju):

10 pecent increase inr so radius=r+10/100*r=11r/10

OpenStudy (anonymous):

A=pi r^2 r/dr =0.1 dA/A=? A= pi *r^2 1) dA/dr =2 pi r 2) dA= 2 pi r dr 3) dividing 3 by 1 dA/A =2r/dr dA/A= 2* 0.1=0.2 20 percent

OpenStudy (anonymous):

@girlish

OpenStudy (anonymous):

listen @amriju it wud be 121% increase isn't it?

OpenStudy (amriju):

no....ur initial area is pi*r^2 write dat as 100/100*pi*r^2....can u..?..now final area is 121pi*r^2/100...so wats the increase..?...21/100*pi*r^2...rite..?...so percent increase=21

OpenStudy (anonymous):

not 21 there might be small error it should be 20 percent

OpenStudy (anonymous):

if u r doing by taking initial values first then there could be some error

OpenStudy (anonymous):

A is prportinal to sqaure of radius that means area will increase by factor of 2

OpenStudy (amriju):

r/dr=10 dear...r is obviously greater than dr...how can u write 0.1

OpenStudy (amriju):

@Fazeelayaz

OpenStudy (anonymous):

well it's 10 percent of the original value 10/100=0.1

OpenStudy (anonymous):

sorr y i have made mistake it's dr/r

OpenStudy (anonymous):

A=pi r^2 dr/r =0.1 dA/A=? A= pi *r^2 1) dA/dr =2 pi r 2) dA= 2 pi r dr 3) dividing 3 by 1 dA/A =2dr/r dA/A= 2* 0.1=0.2 20 percent

OpenStudy (anonymous):

now it's correct

OpenStudy (anonymous):

21 percent is not wrong u took initial values of r that's fine ok but there would be some error in it since pi is irrational u can't wright whole value of it down

OpenStudy (anonymous):

do get me @amriju

OpenStudy (amriju):

no...coz..pi is going to cancel...even if took pi=12147.2147....the answer wud not hav varied...

OpenStudy (amriju):

*even if I

OpenStudy (anonymous):

ok

OpenStudy (amriju):

:)

OpenStudy (anonymous):

now lets take r = 1 ok A=pi r^2 A=pi after increasing by 10 percent area is A2= pi(1.1)^2=pi*1.21 A2-A1=0.21 ok

OpenStudy (anonymous):

it's 21 percent take it that's not wrong but correct way is to do by differentiation

OpenStudy (amriju):

no....differentiation is the process of studing rate of change...not change itself...in any case...no method is wrong as long as it works for any genaralised value...

OpenStudy (anonymous):

i think u should get a medal

OpenStudy (amriju):

oh...thnks....